HW-Time-Slice Firmware-Filter Software vx2740 Steering Module Reports MVM MVM Vexos MVM-Bug listing MVM TRIUMF Local DS Prototype DS Cryogenic For Shifters BCIT-31 ChronoBox
  CERN DS-Proto0 read-only backup, Page 1 of 2  Not logged in ELOG logo
ID Date Author Type Category Subject Text Attachments
  24   18 Nov 2019 02:04 Yi WangProcedure/OperationCooling SystemOpen dewar refill9:00 am open dewar refilled to full, TPC
pressure 30 mbar
  
  23   17 Nov 2019 00:45 Yi WangProcedure/OperationCooling Systemopen dewar refill09:45 am open dewar refill to full, TPC
pressure 31 mbar
  
  22   16 Nov 2019 00:12 Yi WangProcedure/OperationCooling SystemGas pocket formation09:12 am MB OFF, LED ON, start forming
gas pocket

10:46 am Gas pocket 7mm thick,
  
  21   15 Nov 2019 23:54 Yi WangProcedure/OperationCooling Systemopen dewar refill08:50 am open dewar refill to full   
  20   15 Nov 2019 00:14 Yi WangProcedure/OperationCooling System Nov 14, 2019

5:00 pm

Stop fast recirculation, fill open
  
  19   14 Nov 2019 00:34 E. SanchezProcedure/OperationGeneral 9:20 TPC pressure  235 mbar recirculation 
pressure 235 mbar. Gas pocket on. LED OFF.
MB ON.  LAr level  50 cm 
  
  18   13 Nov 2019 05:32 Xiang XiaoProcedure/OperationCooling SystemAfternoon system monitoring14:31 TPC pressure 317 mbar,
recirculation pressure 276 mbar

15:00 TPC pressure 328 mbar,
  
  17   13 Nov 2019 00:26 Yi WangProcedure/OperationCooling SystemMorning system monitoring09:30 Heat Gun ON

09:40 TPC pressure 293 mbar, recirculation
pressure 277 mbar
  
  16   12 Nov 2019 05:47 Edgar SanchezProcedure/OperationGeneralAfternoon system monitoring 14:47 TPC pressure 330 mbar Recirculation
pressure: 290 mbar 

15:35 TPC pressure 345 mbar
  
  15   12 Nov 2019 00:48 Yi WangProcedure/OperationCooling Systemmorning system monitoring09:47 Heat gun ON, TPC pressure 261 mbar,
recirculation pressure 252 mbar

11:00 Building 182 liquid argon
  
  14   11 Nov 2019 10:15 Sam HillProcedure/OperationCooling SystemAfternoon system monitoringChecked pressure regularly, steadily
climbed from 15:00-19:00 

19:00: TPC:
  
  13   11 Nov 2019 04:08 Yi WangProblemCooling SystemPossible power blackout on Nov 9There was a possible power blackout on
Nov 9 from 2 am to 4 am at builidng 182.

During the blackout, the vacuum
  
  12   11 Nov 2019 00:39 Yi WangProcedure/OperationCooling SystemMorning system monitoring09:39 heat gun on

12:37 TPC pressure 350 mbar, recirculation
pressure 338 mbar
  
  11   10 Nov 2019 05:34 Edgar SanchezProcedure/OperationCooling SystemAfternoon system monitoring 14:34 TPC pressure 300 mbar , recirculation
pressure 287 mbar 

15:00 Heaters off for S1 data taking 
  
  10   10 Nov 2019 00:25 Yi WangProcedure/OperationCooling SystemMorning system monitoring09:25 Heat gun ON

09:40 TPC pressure 241 mbar, recirculation
pressure 231 mbar
  
  9   09 Nov 2019 06:35 Sam HillProcedure/OperationCooling SystemAfternoon system monitoring15:36 TPC 320 mbar Recirculation 320 mbar

16:01 TPC 313 mbar Recirculation
313 mbar
  
  8   09 Nov 2019 01:37 Xiang XiaoProcedure/OperationCooling SystemMorning system monitoring10:20 Heater gun on

10:20 TPC 230 mbar Recirculation
241 mbar
  
  7   08 Nov 2019 05:39 Edgar SanchezProcedure/OperationCooling SystemAfternoon system monitoring 14:00 TPC 330 mbar Recirculation 290 mbar.
Liquid argon level in bath 65 cm 

15:00 TPC 335 mbar Recirculation
  
  6   08 Nov 2019 00:28 Yi WangProcedure/OperationCooling SystemMorning system monitoring09:29 Heat Gun ON

09:30 TPC 259 mbar Recirculation
257 mbar
  
  5   07 Nov 2019 05:19 Yi WangProcedure/OperationCooling Systemfast recirculation08:40, switch on heat gun, open dewar liquid
level 50/120 (cm)

16:30, open dewar is refilled to 70/120
  
ELOG V3.1.4-c76272a